a parking meter accepts only quarters and dollars.there are 31 coins value of 20.50 how many quarters and dollars are there
Answer by Alan3354(6097)

a parking meter accepts only quarters and dollars.there are 31 coins value of 20.50 how many quarters and dollars are there
---------------------
Q + D = 31 (# of coins)
25Q + 100D = 2050 (value of coins)
D = 31-Q
-------------
25Q + 100D = 2050
25Q + 100*(31-Q) = 2050
25Q + 3100 - 100Q = 2050
-75Q = -1050
Q = 14
D = 17